# Deep Neural Networks and Brain Alignment: Brain Encoding and Decoding (Survey)

### v2

@@ 3 Naturalistic Neuroscience Datasets
Video datasets. Recently, video neuroscience datasets have also been proposed. These include BBC’s Doctor Who (Seeliger et al., 2019), Japanese Ads (Nishida et al., 2020), Pippi Langkous (Anderson et al., 2020) and Algonauts (Cichy et al., 2021). Japanese Ads data contains data for two sets of movies provided by NTT DATA Corp: web and TV ads. There are also four types of cognitive labels associated with the movie datasets: scene descriptions, impression ratings, ad effectiveness indices, and ad preference votes. Algonauts 2021 contains fMRIs from 10 human subjects that watched over 1,000 short (3 sec) video clips.
@@
Other multimodal datasets. Finally, beyond the video datasets, datasets have also been proposed with other kinds of multimodality. These datasets are audiovisual ((Zinszer et al., 2018; Cao et al., 2021)), words associated with line drawings (Mitchell et al., 2008; Sudre et al., 2012), pictures along with sentences and word clouds (Pereira et al., 2018). These datasets have been collected using a variety of methods like fMRIs (Mitchell et al., 2008; Pereira et al., 2018), MEG (Sudre et al., 2012) and fNIRS (Zinszer et al., 2018; Cao et al., 2021). Specifically, in Sudre et al. (2012), subjects were asked to perform a question answering (QA) task, while their brain activity was recorded using MEG. Subjects were first presented with a question (e.g., “Is it manmade?”), followed by 60 concrete nouns, along with their line drawings, in a random order. For all other datasets, subjects performed passive viewing and/or listening.
Figure 7: Non-invasive brain recordings: fMRI and MEG. This figure is adapted from (Toneva et al., 2022a).
Figure 8: Contrast of estimated cross-subject prediction accuracy for reading and listening for a representative subject (subject-5). Blue and Orange voxels depict higher cross-subject prediction accuracy estimates during listening and reading, respectively. Voxels that have similar cross-subject prediction accuracy during reading and listening appear white, and are distributed across language regions. This Figure is adapted from Oota et al. (2024a).
@@ 4 Brain Regions
In this section, we discuss the mapping of brain recordings to stimulus-specific brain regions that have been discussed in the literature. Specifically, we discuss the regions of language network, auditory cortex and visual cortex, along with their sub regions. To use brain recordings from preprocessed naturalistic neuroscience datasets, follow these steps: (i) Use brain activation of voxels directly if either of the next two steps is applicable. (ii) Apply a brain mask to the brain volume to obtain the activation of voxels. or (iii) Project the brain volume onto the surface space (such as "fsaverage5," "fsaverage6," or "fsaverage"). To visualize the brain maps, popular libraries such as Nilearn  [^2] https://nilearn.github.io/stable/index.html or Pycortex  [^3] https://gallantlab.org/pycortex/ are useful for fMRI recordings, while MNE-Python  [^4] https://mne.tools/stable/index.html is suitable for both MEG and EEG datasets.
@@
Language network. The language network refers to brain regions that are involved in language processing. Based on the Fedorenko lab’s language parcels (Fedorenko et al., 2010; Fedorenko & Thompson-Schill, 2014), eight language-relevant regions encompass broader language regions: angular gyrus (AG), anterior temporal lobe (ATL), posterior temporal lobe (PTL), inferior frontal gyrus (IFG), inferior frontal gyrus orbital (IFGOrb), middle frontal gyrus (MFG), posterior cingulate cortex (PCC) and dorsal medium prefrontal cortex (dmPFC), as shown in Figure 7 (left). These eight language networks are used in several recent studies (Toneva & Wehbe, 2019; Toneva et al., 2022a; Aw & Toneva, 2023; Oota et al., 2024b)Oota et al. (2024b; 2023d; 2023a); Dong & Toneva (2023).
@@
To map brain activations to these eight language regions, prior studies use the multimodal parcellation of the human cerebral cortex based on the Glasser Atlas (which consists of 180 regions of interest in each hemisphere) to report the ROI (region of interest) analysis for the brain maps (Glasser et al., 2016). Overall, the data covers eight language brain ROIs with the following subdivisions: (i) AG: PFm, PGs, PGi, TPOJ2, and TPOJ3; (ii) ATL: STSda, STSva, STGa, TE1a, TE2a, TGv, and TGd; (iii) PTL: A5, STSdp, STSvp, PSL, STV, TPOJ1; (iv) IFG: 44, 45, IFJa, IFSp; (v) MFG: 55b; (vi) IFGOrb: a47r, p47r, a9-46v, (vii) PCC: 31pv, 31pd, PCV, 7m, 23, RSC; and (viii) dmPFC: 9m, 10d, d32.
@@
Figure 8 displays the cross-subject prediction accuracy for reading and listening for a representative sample subject. It illustrates that irrespective of text-evoked or speech-evoked brain activity, high-level information processing occurs in the language regions (indicated by white voxels).
@@
Auditory cortex. The auditory cortex (AC) is a specific brain region responsible for processing auditory information, including the perception of sound, speech, music, and other auditory stimuli. Figure 8 displays the cross-subject prediction accuracy for reading and listening for a representative sample subject. It illustrates that during speech-evoked brain activity, the early auditory cortex (EAC) has higher prediction accuracy (indicated by Blue voxels), signifying early sensory information processing, while high-level information processing occurs in the language regions (indicated by white voxels). Overall, the auditory cortex is divided into the following subdivisions (Nastase et al., 2020): (i) EAC (early auditory cortex): A1 (Primary Auditory Cortex), the Lateral Belt (LBelt), Posterior Belt (PBelt), Medial Belt (MBelt), Rostral Intermediate (RI), and (ii) AAC (auditory association cortex): A4 and A5.
@@
Together, these distinct areas work in concert to form a sophisticated system for perceiving and interpreting the diverse aspects of auditory stimuli.
@@
Visual cortex. The visual cortex is a critical part of the brain responsible for visual information processing, allowing us to see and understand the world around us. Many experiments contrast brain activity elicited by specific image categories. This functional localizer approach has been used to identify many regions of interest (ROIs) in the visual pathway representing information from low-level visual (early) to high-level semantic information. The early visual cortex (EVC) is primarily responsible for processing basic visual information, including detecting of simple features like edges, colors, shapes, and motion. Higher visual cortex (HVC) regions are involved in more advanced visual processing tasks, such as object recognition, facial recognition, scene perception, and the integration of complex visual information.
@@
Overall, the visual cortex is divided into the following subdivisions: (i) PVC (primary visual cortex): V1, (ii) EVC (early visual cortex): V2, V3 and V4, (ii) VWFA (visual word form area) (ii) HVC (high-level visual cortex): the extrastriate body area (EBA), occipital face area (OFA), and the fusiform face area (FFA), the occipital place area (OPA), the parahippocampal place area (PPA), and the retrosplenial cortex (RSC).
@@ 5 Evaluation Metrics
In this section, we discuss popular metrics for evaluation of brain encoding and decoding models.
@@ 5.1 Metrics for Brain Encoding Models
Two metrics are popularly used to evaluate brain encoding models: 2V2 accuracy (Toneva et al., 2020; Oota et al., 2022b) and Pearson Correlation (Jain & Huth, 2018). They are defined as follows.
Given a subject and a brain region, let $N$ be the number of samples. Let $\{Y_{i}\}_{i=1}^{N}$ and $\{\hat{Y}_{i}\}_{i=1}^{N}$ denote the actual and predicted voxel value vectors for the $i^{th}$ sample. Thus, $Y\in R^{N\times V}$ and $\hat{Y}\in R^{N\times V}$ where $V$ is the number of voxels in that region.
@@
2V2 classification accuracy. This metric evaluates how close the brain activity prediction is from ground truth, such as Euclidean distance and cosine distance. This metric evaluates the fMRI predictions using them in a classification task on held-out data in the cross-validation setting. The classification task is to try to match the predicted left-out brain responses to their corresponding ground truth, as introduced in  (Mitchell et al., 2008; Wehbe et al., 2014; Toneva et al., 2020; Aw & Toneva, 2023). Having two sets of brain predictions $\hat{Y}_{i}$ and $\hat{Y}_{j}$, and corresponding ground truth $Y_{i}$ and $Y_{j}$, the 2V2 classification accuracy is computed as $\frac{1}{N_{C_{2}}}\sum_{i=1}^{N-1}\sum_{j=i+1}^{N}I[\{\text{cosD}(Y_{i},\hat{Y}_{i})+\text{cosD}(Y_{j},\hat{Y}_{j})\}<\{\text{cosD}(Y_{i},\hat{Y}_{j})+\text{cosD}(Y_{j},\hat{Y}_{i})\}]$ where cosD is the cosine distance function. $I[c]$ is an indicator function such that $I[c]=1$ if $c$ is true, else it is 0. The higher the 2V2 accuracy, the better. Figure 9 (left) illustrates the computation of 2V2 Accuracy for the case where sample $i$ and $j$ correspond to the brain activity of concepts “dog” and “house”, respectively. This metric was proposed to boost the signal-to-noise ratio in estimating the brain alignment for single-trial data (Aw & Toneva, 2023). Under this metric, chance performance is 50%.
@@
Pearson correlation. This metric evaluates the similarity between the fMRI predictions ($\hat{Y}_{i}$) and the corresponding true fMRI data ($Y_{i}$) by computing the Pearson correlation for each voxel i. The Pearson correlation for voxel i is computed as PCi=$corr[Y_{i},\hat{Y}_{i}]$ where corr is the correlation function. The average Pearson correlation across all voxels is then computed as PCC=$\frac{1}{N}\sum_{i=1}^{n}corr[Y_{i},\hat{Y}_{i}]$, where N denotes the number of voxels. This metric is widely used in cognitive neuroscience (Jain & Huth, 2018; Toneva & Wehbe, 2019; Caucheteux et al., 2021; Goldstein et al., 2022; Aw & Toneva, 2023; Oota et al., 2022b; 2024b).
@@
Cross-subject prediction accuracy. To account for the intrinsic noise in biological measurements and obtain a more accurate estimate of the model’s performance,  Schrimpf et al. (2021) proposed an approach to estimate the cross-subject prediction accuracy. This is achieved by estimating the amount of brain response in one subject that can be predicted using only the data from a combination of other subjects using an encoding model. For instance, consider Harry Potter dataset with n=8 participants, the first step is to subsample—the data with $n$ participants into all possible combinations of $s$ participants for all $s\in[2,8]$ (e.g. 2, 3, 4, 5, 6, 7, 8 for $n$=8). In the second step, for each subsample, select a random participant as the target that we attempt to predict from the remaining $s-1$ participants (e.g., predict 1 subject from 1 (other) subject, 1 from 2 subjects, …, 1 from 8, to obtain a mean score for each voxel in that subsample. In the third step, extrapolate to infinitely many humans and thus to obtain the highest possible (most conservative) estimate, as suggested by Schrimpf et al. (2021), fit the equation $v=v_{0}\times\left(1-e^{-\frac{x}{\tau_{0}}}\right)$ where $x$ is each subsample’s number of participants, $v$ is each subsample’s correlation score and $v_{0}$ and $\tau_{0}$ are the fitted parameters. This fitting was performed for each voxel independently with 100 bootstraps each to estimate the variance where each bootstrap draws $x$ and $v$ with replacement. The final ceiling value was the median of the per-voxel ceilings $v_{0}$.
@@
Normalized brain alignment. The neural model predictivity values were normalized by their respective subject estimated cross-subject prediction accuracies, as proposed by Schrimpf et al. (2021). The final measure of a model’s performance (‘normalized brain alignment’ or ‘score’) on a dataset is thus Pearson’s correlation between model predictions and neural recordings divided by the estimated ceiling and averaged across voxel locations and participants.
Figure 9: Evaluation Metrics for Brain Encoding and Decoding. (Left) 2V2 Accuracy (Toneva et al., 2020), (Right) Pairwise Accuracy

### v3

@@ 3.2 Vision datasets
Video datasets. Recently, video neuroscience datasets have also been proposed. These include BBC’s Doctor Who (Seeliger et al., 2019), Japanese Ads (Nishida et al., 2020), Pippi Langkous (Anderson et al., 2020), and Algonauts (Cichy et al., 2021). The Japanese Ads dataset contains data for two sets of movies provided by NTT DATA Corp: web and TV ads. It also includes four types of cognitive labels associated with the movie datasets: scene descriptions, impression ratings, ad effectiveness indices, and ad preference votes. The Algonauts 2021 dataset contains fMRIs from 10 human subjects who watched over 1,000 short (3-second) video clips.
3.3 Multimodal datasets.
@@
Beyond audio and vision datasets, other multimodal datasets have been proposed. These include audiovisual datasets (Zinszer et al., 2018; Cao et al., 2021; Boyle et al., 2020; Jung et al., 2024), and words associated with line drawings (Mitchell et al., 2008; Sudre et al., 2012). These datasets have been collected using various methods such as fMRI (Mitchell et al., 2008; Pereira et al., 2018), MEG (Sudre et al., 2012), and fNIRS (Zinszer et al., 2018; Cao et al., 2021). Specifically, in Sudre et al. (2012), subjects were asked to perform a question answering (QA) task while their brain activity was recorded using MEG. Subjects were first presented with a question (e.g., “Is it manmade?”), followed by 60 concrete nouns along with their line drawings, in a random order. For all other datasets, subjects performed passive viewing and/or listening tasks. The recent multimodal brain datasets, Neuromod (Boyle et al., 2020) and Multimodal fMRI (Jung et al., 2024), involve experiments with multimodal naturalistic stimuli, where participants engage in watching videos accompanied by audio.
Figure 6: Non-invasive Brain Recording Techniques: fMRI and MEG. The fMRI plot illustrates the language network, highlighting the brain regions involved in language processing. In contrast, the MEG plot demonstrates the timing of peak responses to words, emphasizing the temporal aspects of neural activity. Figure is adapted from (Toneva et al., 2022a), used with permission from the respective authors.
@@ 4 Brain Regions
This section discusses the mapping of brain recordings to stimulus-specific brain regions as presented in the literature. We focus on the regions of the language network, auditory cortex, and visual cortex, along with their sub-regions. To use brain recordings from preprocessed naturalistic neuroscience datasets, follow these steps: (i) Use brain activation of voxels directly if either of the next two steps is not applicable. (ii) Apply a brain mask to the brain volume to obtain the activation of voxels. or (iii) Project the brain volume onto the surface space (such as “fsaverage5”, “fsaverage6”, or “fsaverage”). To visualize the brain maps, popular libraries such as Nilearn  [^2] https://nilearn.github.io/stable/index.html or Pycortex  [^3] https://gallantlab.org/pycortex/ are useful for fMRI recordings, while MNE-Python  [^4] https://mne.tools/stable/index.html is suitable for both MEG and EEG datasets.
@@
Language network. The language network refers to brain regions involved in language processing. Studies such as those by Fedorenko & Thompson-Schill (2014); Lipkin et al. (2022) use contrasts like words versus non-words to functionally select language-specific voxels, highlighting the variability of anatomical regions across individuals. As a result, the language network is typically defined functionally rather than anatomically. Based on the Fedorenko lab’s language parcels (Fedorenko et al., 2010; Fedorenko & Thompson-Schill, 2014), eight language-relevant regions encompass broader language areas: angular gyrus (AG), anterior temporal lobe (ATL), posterior temporal lobe (PTL), inferior frontal gyrus (IFG), inferior frontal gyrus orbital (IFGOrb), middle frontal gyrus (MFG), posterior cingulate cortex (PCC), and dorsal medial prefrontal cortex (dmPFC), as shown in Figure 6 (left). These eight language networks are used in several recent studies (Toneva & Wehbe, 2019; Toneva et al., 2022a; Aw & Toneva, 2023; Oota et al., 2024b; 2023d; 2023a; Dong & Toneva, 2023).
@@
To map brain activations to these eight language regions, prior studies use the multimodal parcellation of the human cerebral cortex based on the Glasser Atlas (Glasser et al., 2016), which consists of 180 regions of interest in each hemisphere. The data covers eight language brain ROIs with the following subdivisions: (i) AG: PFm, PGs, PGi, TPOJ2, and TPOJ3; (ii) ATL: STSda, STSva, STGa, TE1a, TE2a, TGv, and TGd; (iii) PTL: A5, STSdp, STSvp, PSL, STV, TPOJ1; (iv) IFG: 44, 45, IFJa, IFSp; (v) MFG: 55b; (vi) IFGOrb: a47r, p47r, a9-46v; (vii) PCC: 31pv, 31pd, PCV, 7m, 23, RSC; and (viii) dmPFC: 9m, 10d, d32.
@@
Figure 7 displays the cross-subject prediction accuracy for reading and listening for a representative sample subject from the subset-moth-radio-hour dataset (Oota et al., 2024a). It illustrates that, irrespective of text-evoked or speech-evoked brain activity, high-level information processing occurs in the language regions (indicated by white voxels).
Figure 7: Contrast of estimated cross-subject prediction accuracy for reading and listening for a representative subject (subject-5 from the subset-moth-radio-hour dataset). Blue and Orange voxels depict higher cross-subject prediction accuracy estimates during listening and reading, respectively. Voxels that have similar cross-subject prediction accuracy during reading and listening appear white, and are distributed across language regions. This figure is adapted from (Oota et al., 2024a), and used with permission from the respective authors.
@@
Auditory cortex. The auditory cortex (AC) is a specific brain region responsible for processing auditory information, including the perception of sound, speech, music, and other auditory stimuli. Figure 7 illustrates that during speech-evoked brain activity, the early auditory cortex (EAC) has higher prediction accuracy (indicated by blue voxels), signifying early sensory information processing, while high-level information processing occurs in the language regions (indicated by white voxels). The auditory cortex is divided into the following subdivisions (Nastase et al., 2020): (i) EAC (early auditory cortex): A1 (Primary Auditory Cortex), the Lateral Belt (LBelt), Posterior Belt (PBelt), Medial Belt (MBelt), Rostral Intermediate (RI), and (ii) AAC (auditory association cortex): A4 and A5.
@@
Together, these distinct areas work in concert to form a sophisticated system for perceiving and interpreting diverse aspects of auditory stimuli.
@@
Visual cortex. The visual cortex is a critical part of the brain responsible for visual information processing, allowing us to see and understand the world around us. Many experiments contrast brain activity elicited by specific image categories. This functional localizer approach has been used to identify many regions of interest (ROIs) in the visual pathways (dorsal and ventral streams), representing information from low-level visual (early) to high-level semantic information.
The early visual cortex (EVC) is primarily responsible for processing basic visual information, including the detection of simple features like edges, colors, shapes, and motion. After initial processing in the early visual cortex, visual information is transmitted along two pathways (dorsal and ventral streams) for higher-level interpretation. The ventral pathway, often referred to as the “what” pathway, extends from the EVC to the inferior temporal cortex (IT). This pathway is crucial for object recognition and form representation. It handles the processing of detailed visual information, enabling us to identify and understand complex stimuli such as faces, objects, and scenes. In contrast, the dorsal pathway, known as the “where” or “how” pathway, projects from the EVC to the posterior parietal cortex. It is primarily responsible for spatial awareness and motion detection. The dorsal stream processes information about the location and movement of objects in space, facilitating the coordination of actions like reaching, grasping, and navigating through the environment. In summary, the Higher visual cortex (HVC) includes regions from both the dorsal and ventral pathways, each specializing in different aspects of visual processing.
@@
Overall, the visual cortex is divided into the following subdivisions: (i) PVC (primary visual cortex): V1, (ii) EVC (early visual cortex): V2, V3 and V4, (iii) VWFA (visual word form area), and (iv) HVC (high-level visual cortex): Ventral visual regions - the extrastriate body area (EBA), occipital face area (OFA), and the fusiform face area (FFA), the occipital place area (OPA), the parahippocampal place area (PPA), and the retrosplenial cortex (RSC); Dorsal visual regions - V3A and V3B (associated with motion processing), V6 and V6A (integrate visual and somatic inputs to inform motor actions), V7 and IPS1 (involved in visual attention and mapping objects in space), and MT & MST (play roles in detecting and analyzing moving objects in our environment).
@@ 5 Evaluation Metrics
This section discusses popular metrics for evaluating brain encoding and decoding models.
@@ 5.1 Metrics for Brain Encoding Models
Given a subject and a brain region, let $N$ be the number of samples. Let ${\text{Y}_{i}}$ and ${\hat{\text{Y}}_{i}}$ denote the actual and predicted voxel value vectors for the $i^{th}$ sample. Thus, $\text{Y}\in\mathbb{R}^{N\times V}$ and $\hat{\text{Y}}\in\mathbb{R}^{N\times V}$, where $V$ is the number of voxels in that region. The process comprises two steps, one estimating the predicted value ($\hat{\text{Y}}$) and then comparing with the ground truth value (Y). Predictions are typically estimated from either linear models such as partial least squares regression (PLS) (Yamins et al., 2014; Schrimpf et al., 2020), ridge regression or other nonlinear models. The predictions from brain encoding models are compared with ground truth values using two common metrics: 2V2 accuracy (Toneva et al., 2020) and Pearson Correlation. They are defined as follows:
@@
2V2 classification accuracy. This metric evaluates how close the brain activity prediction is to ground truth, using measures such as Euclidean distance or cosine similarity. It assesses fMRI predictions by using them in a classification task on held-out data in a cross-validation setting. The task is to match predicted left-out brain responses to their corresponding ground truth, as introduced in (Mitchell et al., 2008; Wehbe et al., 2014; Toneva et al., 2020; Aw & Toneva, 2023). Given two sets of brain predictions $\hat{\text{Y}}_{i}$ and $\hat{\text{Y}}_{j}$, and corresponding ground truth $\text{Y}_{i}$ and $\text{Y}_{j}$, the 2V2 classification accuracy is computed as:
$\frac{1}{\binom{N}{2}}\sum_{i=1}^{N-1}\sum_{j=i+1}^{N}\mathbb{I}[\cos(\text{Y}_{i},\hat{\text{Y}}_{i})+\cos(\text{Y}_{j},\hat{\text{Y}}_{j})>\cos(\text{Y}_{i},\hat{\text{Y}}_{j})+\cos(\text{Y}_{j},\hat{\text{Y}}_{i})]$ (1)
where $\cos(\text{A},\text{B})$ denotes the cosine similarity between vectors A and B, and $\mathbb{I}[c]$ is an indicator function such that $\mathbb{I}[c]=1$ if $c$ is true, else it is 0. Higher 2V2 accuracy indicates better performance. Figure 8 (left) illustrates the computation of 2V2 Accuracy for the case where samples $i$ and $j$ correspond to the brain activity of concepts “dog” and “house”, respectively. This metric was proposed to boost the signal-to-noise ratio in estimating brain alignment for single-trial data (Aw & Toneva, 2023). Under this metric, chance performance is 50%.
@@
Pearson correlation. This metric evaluates the similarity between fMRI predictions ($\hat{\text{Y}}_{i}$) and the corresponding true fMRI data ($\text{Y}_{i}$) by computing the Pearson correlation for each voxel $i$. The Pearson correlation for voxel $i$ is computed as:
$\text{PC}_{i}=\text{corr}[\text{Y}_{i},\hat{\text{Y}}_{i}]$ (2)
where corr is the correlation function. The average Pearson correlation across all voxels, Pearson Correlation Coefficient (PCC), is then computed as:
$\text{PCC}=\frac{1}{V}\sum_{i=1}^{V}\text{corr}[\text{Y}_{i},\hat{\text{Y}}_{i}]$ (3)
where $V$ denotes the number of voxels. This metric is widely used in cognitive neuroscience (Jain & Huth, 2018; Toneva & Wehbe, 2019; Caucheteux et al., 2021a; Goldstein et al., 2022; Aw & Toneva, 2023; Oota et al., 2022b; 2024b).
Figure 8: Evaluation Metrics for Brain Encoding and Decoding. (Left) Illustration of 2V2 Accuracy computation, a metric used in brain encoding models to evaluate prediction accuracy. (Right) Demonstration of Pairwise Accuracy calculation, commonly used in brain decoding models. This figure is from Pereira et al. (2018), and used with permission from the authors.
@@
Cross-subject prediction accuracy. To account for intrinsic noise in biological measurements and obtain a more accurate estimate of model performance, Schrimpf et al. (2021) proposed an approach to estimate cross-subject prediction accuracy. This is achieved by estimating the amount of brain response in one subject that can be predicted using only data from a combination of other subjects using an encoding model. For instance, consider the Harry Potter dataset with $n=8$ participants. The process begins by subsampling the data with $n$ participants into all possible combinations of $s$ participants for all $s\in[2,8]$ (e.g., 2, 3, 4, 5, 6, 7, 8 for $n=8$). For each of these subsamples, a random participant is selected as the target to be predicted from the remaining $s-1$ participants. This approach allows for predictions ranging from 1 subject based on 1 other subject, to 1 subject based on 2 subjects, and so on up to 1 subject based on 7 subjects, ultimately yielding a mean score for each voxel in that subsample. To extrapolate these results to infinitely many humans and obtain the highest possible (most conservative) estimate, the following equation is fitted:
$v=v_{0}\times\left(1-e^{-\frac{x}{\tau_{0}}}\right)$ (4)
In this equation, $x$ represents each subsample’s number of participants, $v$ denotes each subsample’s correlation score, and $v_{0}$ and $\tau_{0}$ are the fitted parameters. This fitting procedure is performed independently for each voxel, using 100 bootstraps to estimate variance. Each bootstrap iteration draws $x$ and $v$ with replacement. The final ceiling value is determined by taking the median of the per-voxel ceilings $v_{0}$. This approach provides a robust estimate of cross-subject prediction accuracy, accounting for the variability inherent in biological measurements.
@@
Normalized brain alignment. The neural model predictivity values can be normalized by their respective subject estimated cross-subject prediction accuracies, as proposed by Schrimpf et al. (2021). Thus, normalized brain alignment on a dataset is computed as Pearson’s correlation between model predictions and neural recordings divided by the estimated ceiling and averaged across voxel locations and participants.